A child throws a 0.50 kg snowball at his friend at 12.0 m/s but the snowball lands on the ground in front of his friend.

The snowball is 1.3 m above the ground when it leaves the child’s hand.


(a) What is the (i) kinetic energy & (ii) potential energy of the snowball at the time of release?


(b) What is the total energy of the snowball (i) at the time of release & (ii) just before it hits the ground?


(c) What is the snowball’s kinetic energy just before it hits the ground in front of his friend?


(d) How fast is the snowball moving just before it hits the ground in front of his friend?



1
Expert's answer
2022-04-20T13:47:09-0400

"K=\frac{mv^2}2=36~J,"

"P=mgh=6.4~J,"

"E=K+P=42.4~J,"

"E_1=E_2=E=42.4~J,"

"v_1=v=12~\frac ms."